# font=small

### v1

@@ 3 A Stochastic Optimal Control Market-Making Problem
A MM problem, in the SOC sense, involves a financial market player who would like to maximize their terminal wealth by trading a large amount of a financial asset, often using many LOs. This type of trader is often referred to as a liquidity provider, which refers to their ability to post a large amount of LOs throughout the LOB. In the example problem we give, the control variable is whether to be posted on the best bid/ask, which is often a decision a MM tries to make. What makes this particular trading problem special is that it begins to incorporate adverse selection in the trading process, which is a big hurdle for a market-maker in reality.
Firstly, we would like to introduce the MM problem we studied, as given in Cartea et al., (2015). We begin by describing the key stochastic processes as follows,
  • $\delta^{\pm}=(\delta_{t}^{\pm})_{\{0\leq t\leq T\}}$, where $\delta_{t}^{\pm}\in\{0,1\}$, is the agents control process, which tells the market-maker whether or not to be posted at the best bid/ask.
  • $Q^{\delta}=(Q_{t}^{\delta})_{\{0\leq t\leq T\}}$ is the agents’ controlled inventory process and is impacted by how much the agent trades.
  • $S^{\delta}=(S_{t}^{\delta})_{\{0\leq t\leq T\}}$ is the controlled midprice process of the financial asset being traded, which is also impacted by how much the agent trades.
@@
    $dC_{t}^{\delta}=\displaystyle{\left(S_{t}+\frac{\Delta}{2}\right)}dN_{t}^{\delta,+}-\displaystyle{\left(S_{t}-\frac{\Delta}{2}\right)}dN_{t}^{\delta,-}$ (4)
@@
where $\mathscr{A}$ is the set of admissible strategies in which $\delta>0$ and uniformly bounded from above. It is also important to note that there is an inventory constraint in this problem, whereby the inventory process has an upper $\overline{q}$ and lower $\underline{q}$ bound. This essentially sets the maximum position the market maker can take, long or short. Then, by applying the Dynamic Programming Principle (DPP), as explained in Cartea et al., (2015), the value function should satisfy the following Dynamic Programming Equation (DPE),
$\begin{split}0=\left(\partial_{t}+\alpha\partial_{s}+\frac{1}{2}\sigma^{2}\partial_{ss}-\zeta\alpha\partial_{\alpha}+\frac{1}{2}\eta^{2}\partial_{\alpha\alpha}\right)H-\phi q^{2}\\ +\lambda^{+}\max_{\delta^{+}\in\{0,1\}}\{\mathbb{I}_{q>\underline{q}}\mathbb{E}\left[H(t,c+\left(S+\frac{\Delta}{2}\delta^{+}\right)\delta^{+},S,\alpha+\epsilon^{+},q-\delta^{+})-H\right]\\ +(1-\delta^{+}\mathbb{I}_{q>\underline{q}})\mathbb{E}\left[H(t,c,S,\alpha+\epsilon^{+},q)-H\right]\}\\ +\lambda^{-}\max_{\delta^{-}\in\{0,1\}}\{\mathbb{I}_{q<\overline{q}}\mathbb{E}\left[H(t,c-\left(S-\frac{\Delta}{2}\delta^{-}\right)\delta^{-},S,\alpha-\epsilon^{-},q+\delta^{-})-H\right]\\ +(1-\delta^{-}\mathbb{I}_{q<\overline{q}})\mathbb{E}[H(t,c,S,\alpha-\epsilon^{-},q)-H]\}\end{split}$ (7)
@@
  • Line 1: Characterizes the drift and diffusive parts of the midprice and short-term alpha processes, along alpha’s mean-reverting component. The last term on this line includes the running inventory penalty component, which only effects the inventory process.
  • Line 2: Characterizes the change in the value function if the agent has a sell LO posted and the effect of an arriving buy MO filling the agent’s posted sell LO. This filled LO creates a jump in the short-term alpha component. The maximisation term characterizes the agent’s control, which determines whether to post a LO at the best offer.
@@
  • Line 4-5: Analogous to lines 2-3 but for the buy side of the LOB.
In Cartea et al., (2015), they use an ansatz to simplify the DPE in equation (7). This ansatz splits out the accumulated cash, the book value of the shares which are marked-to-market at the midprice and the added value from making markets optimally as follows:
@@
$\begin{split}0=\displaystyle\left(\partial_{t}-\zeta\alpha\partial_{\alpha}+\frac{1}{2}\eta^{2}\partial_{\alpha\alpha}\right)h+\alpha q-\phi q^{2}\\ +\lambda^{+}\max_{\delta^{+}\in\{0,1\}}\displaystyle\left\{\mathbb{I}_{q>\underline{q}}\mathbb{E}\displaystyle\left[\delta^{+}\frac{\Delta}{2}+h(t,\alpha+\epsilon^{+},q-\delta^{+})-h(t,\alpha+\epsilon^{+},q)\right]\right\}\\ +\lambda^{-}\max_{\delta^{-}\in\{0,1\}}\displaystyle\left\{\mathbb{I}_{q<\overline{q}}\mathbb{E}\displaystyle\left[\delta^{-}\frac{\Delta}{2}+h(t,\alpha-\epsilon^{-},q+\delta^{-})-h(t,\alpha-\epsilon^{-},q)\right]\right\}\\ +\lambda^{+}\mathbb{E}[h(t,\alpha+\epsilon^{+},q)-h(t,\alpha,q)]\\ +\lambda^{-}\mathbb{E}[h(t,\alpha-\epsilon^{-},q)-h(t,\alpha,q)]\end{split}$ (10)
subject to the terminal condition,
@@
Lastly, by letting $\delta=0$, the optimal postings can be represented in a compact form due to the structure of the maximization terms as follows,
$\begin{split}\delta^{+,*}(t,\alpha,q)=\mathbb{I}_{\{\frac{\Delta}{2}+\mathbb{E}[h(t,\alpha+\epsilon^{+},q-1)-h(t,\alpha+\epsilon^{+},q)]>0\}\cap\{q>\underline{q}\}}\\ \delta^{-,*}(t,\alpha,q)=\mathbb{I}_{\{\frac{\Delta}{2}+\mathbb{E}[h(t,\alpha-\epsilon^{-},q+1)-h(t,\alpha-\epsilon^{-},q)]>0\}\cap\{q<\overline{q}\}}\\ \end{split}$ (12)
Here, we can see that the decision to post depends on the half-spread plus the expectation over the value functions computed at $\alpha\pm\epsilon^{\pm}$, where an arriving MO leads to a positive/negative jump in $\alpha$.
@@ 4 Trading Simulation Environment
In this section we discuss the trading simulation environment used to test the performance of the optimal MM strategy. In Cartea et al., (2015) and Jaimungal, (2019), they create a simulation environment in order to back-test this strategy, which we used as our benchmark. We will make one update to the benchmark by using real LOB data, rather than simulated prices. This enabled us to back-test the strategy on real financial futures contracts data later on. We will discuss the extensions we made in our improved simulation environment, which we believe more accurately portrays how this type of MM strategy would perform in reality, albeit, still with some limitations. Specifically, we will describe how our framework tracks the adverse fills, as well as an improved fill probability structure related to non-adverse fills. We will then proceed to show how this can have a major effect on the performance of the trading strategy.
@@
  • Asset price: For this purpose, we will use Futures LOB data, which was collected at the proprietary trading firm Futures First Inc. The products we will focus on include the near-term expiry contracts for ES, NQ, CL and ZN. Market makers often tend to focus on the more liquid products like these, in particular for HFT, as it will enable them to trade more volume. We resampled our price data into a 1 second time interval, which was then subsequently used as the time step size in our discrete environment.
@@
  • Trade order fills: The benchmark trading simulation environment essentially assumes all fills are non-adverse, unless in the unlikely event a fill occurs at the exact same time as an adverse price movement. In our improved environment, however, we include adverse fills, as well as non-adverse fills, in the fill tracking process. The benchmark environment also assumes that all trade orders go to the front of the LOB queue, thus, their LOs will be the first to be filled. In most liquid financial assets, this is not very likely to occur. Based on this and using the empirical evidence from Tables 1 and 2 earlier, we will attach a binomial probability $p$ to all non-adverse fills. And so, non-adverse fills can be defined as follows,
    $NFA_{t}=\sum_{i=1}^{N}\delta^{+}_{t_{i}}\mathbb{I}_{\{M^{+}_{t_{i}}=1\}}*p$ (13)
    and
    $NFB_{t}=\sum_{i=1}^{N}\delta^{-}_{t_{i}}\mathbb{I}_{\{M^{-}_{t_{i}}=1\}}*p,$ (14)
    for $i=0,...,N$, where $N$ is the total number of time steps. Here, $NFA_{t}$ and $NFB_{t}$ are the counting processes for all non-adverse fills on the best ask and best bid, respectively. $p$ can be defined as:
    $p=P(N_{t_{i}}^{\delta,\pm}|\delta_{t_{i}}^{+}\mathbb{I}_{\{AS(t_{i})\geq AS(t_{i+1})\}}=0,\delta_{t_{i}}^{-}\mathbb{I}_{\{BS(t_{i})\leq BS(t_{i+1})\}}=0)$ (15)
    which is the probability of getting filled given the fill is non adverse. In Cartea et al., (2015) and Jaimungal, (2019), although not explicitly defined, this probability would be one, i.e., $p=1$, and thus we keep it like this in the benchmark environment. Next, we will keep track of the adverse fills in the following manner:
    $AFA_{t}=\sum_{i=1}^{N}\delta_{t_{i}}^{+}\mathbb{I}_{\{AS(t_{i})<AS(t_{i+1})\}},$ (16)
    and
    $AFB_{t}=\sum_{i=1}^{N}\delta_{t_{i}}^{-}\mathbb{I}_{\{BS(t_{i})>BS(t_{i+1})\}},$ (17)
@@
    for $i=0,...,N$, where $AFA_{t}$ and $AFB_{t}$ are counting processes for adverse fills on the ask and adverse fills on the bid, respectively. Here, $BS_{t_{i}}$ and $AS_{t_{i}}$ represent the bid and ask prices, respectively. To understand this intuitively, at each point in time, the trading strategy has a LO posted on the best bid (ask) $\delta_{t_{i}}^{-}(\delta_{t_{i}}^{+})$ and the asset price at time step $t_{i+1}$ is lower (higher) than at time $t_{i}$, the traders LO was then filled at the price $BS_{t_{i}}(AS_{t_{i}})$. This is quite logical because, in reality, based on the rules of the LOB system, the asset price can not move below (above) a traders posted bid (ask) LO without filling their LO first. All other fills the trader receives will be considered non-adverse fills. Finally, in order to determine whether a fill occurred or not, we combine the two fill tracking formulas as follows,
    $N_{t}^{\delta,+}=\sum_{i=1}^{N}max(AFA_{t_{i}},NFA_{t_{i}})$ (18)
    and
    $N_{t}^{\delta,-}=\sum_{i=1}^{N}max(AFB_{t_{i}},NFB_{t_{i}})$ (19)
@@
    where $N_{t}^{\delta,+}$ and $N_{t}^{\delta,-}$ are now counting processes collecting all the trade order fills that occur throughout this particular trading strategy.
@@ 5 Results
Here, we will discuss the results from testing the SOC MM strategy in our improved simulation environment and we will compare this to the benchmark environment. First, we would like to briefly summarize the numerical solution under our parameter set and we provide a visual in Figure 2, which looks similar to the solution given in Cartea et al., (2015) and Jaimungal, (2019). One can see how the optimal solution varies as a function of time ($t$), short-term alpha ($\alpha$) and inventory($q$). The left (right) Figure portrays a visual of LOs posted on the best offer (bid), where a sell (buy) LO is posted if the MMs inventory is above (below) the surface. The images look almost identical as the arrival rates of MOs are the same on both sides of the LOB. Far from maturity, the strategy is mostly independent of time, but one can see that as the MM approaches the maturity time ($T$), the optimal strategy is essentially independent of the short-term alpha ($\alpha$), since the MM is solely focused on liquidating any remaining position.
Figure 2: Optimal solution as a function of time, short-term alpha and inventory.
@@
As discussed earlier, the main difference in our improved simulation environment is that adverse fills are now included in the strategy simulation process, determined based on equations (16)-(17). Our aim in this section is to show how adverse fills and fill probabilities on non-adverse fills can significantly effect the performance of the trading strategy. We will discuss the results for CL in depth here, where the same repeated analysis for ES, NQ and ZN can be found in the appendix. Based on the empirical evidence for the distribution between adverse and non-adverse fills given in Table 2, we also set the non-adverse fill probability to $p=0.2$. See Table 3, below, for a list of the parameter values used. We kept most of the parameters similar to Cartea et al., (2015) and Jaimungal, (2019), however, we decreased the time step and increased the maturity time, in order to make the length of time more realistic to the data. We also decreased the intensity of the MOs as another way to bring the distribution of adverse and non-adverse fills a bit closer to reality. A few slight adjustments to some of the other parameters were also needed to keep the numerical solution stable.
@@
Now, we will review the strategy simulation from our improved simulation environment alongside the performance of the benchmark environment. In Figure 3, we begin by showing a snapshot of the strategy over a random 120 second path in CL, where one can see when the market-maker is posted on the best bid/ask and when they receive trade order fills. The left and right Figures show the strategy simulation in the benchmark and improved environments, respectively. Here, the green (blue) lines indicate when the market-maker is posted on the best bid (ask), the filled circles indicate when the market-makers LO is filled on the best bid/ask, and the unfilled circles indicate times the market-maker would have been filled if they had a LO posted. One noticeable feature, in the left Figure, is that whenever the agent is posted and price moves through their order, they do not automatically receive a fill. This, as we mentioned earlier, is because MOs are simulated independently from the price process in Cartea et al., (2015), Cartea et al., 2018b and Jaimungal, (2019), and is contrary to what would happen in reality. And so, all the fills in this left Figure would be referred to as non-adverse fills. In the right Figure, one can see where all the adverse fills would have occurred (denoted by AFB and AFA). Note, this can also alter the posting strategy later on because the inventory process now evolves differently to how it would in the benchmark environment.
Figure 3: A snapshot of a random path of the strategy in the benchmark (left) and improved (right) simulation environments. The top line (black/blue) indicates the best ask, the middle line (black) the midprice and the bottom line (black/green) the best bid. When the top (bottom) line is blue (green), the agent is posted on the best ask (bid). Closed circles indicate fills received when the market-maker is posted and open circles indicate fills that the market-maker would’ve received if they were posted.
It’s quite apparent that the timing of the trade order fills has changed and many trade order fills occur at very unfavourable prices. Earlier, in our simple trading example, where we gave results in Table 2, we saw there that a overwhelming majority of the fills for the strategy simulation in CL were adverse, as well as for ES, NQ and ZN. This test was also conducted on the exact same trading day as this simulation. In Table 4, we show the number of each type of fill that occurred within our improved simulation environment. We can see that using a much lower non-adverse fill probability of $p=0.2$ created a similar distribution of fills between non-adverse fills and adverse fills as in Table 2, and we will soon see that the overall performance has now dropped significantly. This trend also persists in the strategy simulations using ES, NQ and ZN data, as can be seen in the appendix. Based on our empirical evidence, we believe these results more accurately portray how the strategy would perform in reality. However, we may not want to be overly pessimistic by thinking the SOC MM strategy should have the same number of adverse fills as the basic posting strategy, because the SOC MM model has a short-term alpha predictor which could potentially predict some of these adverse fills. However, it’s extremely unlikely that any short-term alpha predictor could foresee all future adverse fills.

### v2

@@ 4 A Stochastic Optimal Control Market-Making Problem
An MM problem, in the SOC sense, involves a financial market player who would like to maximize their terminal wealth by trading a large amount of a financial asset, often using many LOs. This type of trader is often referred to as a liquidity provider, which refers to their ability to post a large amount of LOs throughout the LOB. In the example problem we give, the control variable is whether to be posted on the best bid/ask, which is often a decision a MM tries to make. What makes this particular trading problem special is that it begins to incorporate adverse selection in the trading process, which is a big hurdle for a MM in reality.
Firstly, we would like to introduce the MM problem studied, as given in Cartea et al., (2015), where we also introduce an additional probability measure for non-adverse fills. We begin by describing the key stochastic processes as follows,
  • $\delta^{\pm}=(\delta_{t}^{\pm})_{0\leq t\leq T}$, where $\delta_{t}^{\pm}\in\{0,1\}$, represents the agent’s control process. This variable determines whether the MM is posted at the best bid/ask. In our setting, each posted unit is of size 1.
  • $\rho^{\delta}=(\rho^{\delta}_{t})_{0\leq t\leq T}$ denotes the probability that a posted limit order will be filled, where the agent has one unit posted at the best bid/ask. As in Cartea et al., (2015) and our reasoning given above, it is assumed that all trade order fills in this problem are non-adverse. However, the improved simulation framework later on will include adverse fills to show how performance can be significantly affected. Due to the dimension reduction that is needed and shown later on, adverse fills cannot be included in the problem setup. We highlight why this shortcoming occurs later.
  • $Q^{\delta}=(Q_{t}^{\delta})_{\{0\leq t\leq T\}}$ is the agents’ controlled inventory process and is impacted by the amount the agent trades.
  • $S^{\delta}=(S_{t}^{\delta})_{\{0\leq t\leq T\}}$ is the controlled midprice process of the financial asset being traded, which is also impacted by the amount the agent trades.
@@
    $dC_{t}^{\delta}=\displaystyle{\left(S_{t}+\frac{\Delta}{2}\right)}dN_{t}^{\delta,+}-\displaystyle{\left(S_{t}-\frac{\Delta}{2}\right)}dN_{t}^{\delta,-}.$ (4)
@@
where $\mathscr{A}$ is the set of admissible strategies in which $\delta>0$ and uniformly bounded from above. It is also important to note that there is an inventory constraint in this problem, by which the inventory process has an upper $\overline{q}$ and lower $\underline{q}$ bound. This essentially sets the maximum position that the MM can take long or short. Then, by applying the Dynamic Programming Principle (DPP), as explained in Cartea et al., (2015), the value function should satisfy the following Dynamic Programming Equation (DPE),
$\begin{split}0&=\left(\partial_{t}+\alpha\partial_{s}+\frac{1}{2}\sigma^{2}\partial_{ss}-\zeta\alpha\partial_{\alpha}+\frac{1}{2}\eta^{2}\partial_{\alpha\alpha}\right)H-\phi q^{2}\\ &\quad+\lambda^{+}\bigg{(}\max_{\delta^{+}\in\{0,1\}}\bigg{\{}\rho(\delta^{+})\mathbb{I}_{q>\underline{q}}\mathbb{E}\Big{[}\delta^{+}\frac{\Delta}{2}+H(t,c+(S+\frac{\Delta}{2}\delta^{+})\delta^{+},S,\alpha+\epsilon^{+},\\ &\qquad q-\delta^{+})-H\Big{]}\bigg{\}}+\mathbb{E}\Big{[}H(t,c,S,\alpha+\epsilon^{+},q)-H\Big{]}\bigg{)}\\ &\quad+\lambda^{-}\bigg{(}\max_{\delta^{-}\in\{0,1\}}\bigg{\{}\rho(\delta^{-})\mathbb{I}_{q<\overline{q}}\mathbb{E}\Big{[}\delta^{-}\frac{\Delta}{2}+H(t,c-(S-\frac{\Delta}{2}\delta^{-})\delta^{-},S,\alpha-\epsilon^{-},\\ &\qquad q+\delta^{-})-H\Big{]}\bigg{\}}+\mathbb{E}\Big{[}H(t,c,S,\alpha-\epsilon^{-},q)-H\Big{]}\bigg{)},\end{split}$ (7)
@@
  • Line 1: Characterizes the drift and diffusive parts of the midprice and short-term alpha processes, along alpha’s mean-reverting component. The last term on this line includes the running inventory penalty component, which only affects the inventory process.
  • Line 2: Characterizes the change in the value function if the agent has a sell LO posted and the effect of an arriving buy MO filling the agent’s posted sell LO, where it is clear that the midprice is unchanged, thus this represents a non-adverse trade order fill. This filled LO creates a jump in the short-term alpha component. The maximization term characterizes the agent’s control, which determines whether to post an LO at the best offer.
@@
  • Lines 4-5: Analogous to lines 2-3 but for the buy side of the LOB.
In Cartea et al., (2015), they use an ansatz to simplify the DPE in Equation (7) to find a tractable solution. This ansatz splits out the accumulated cash, the book value of the shares which are marked-to-market at the midprice, and the added value from making markets optimally as follows:
@@
$\begin{split}0&=\left(\partial_{t}-\zeta\alpha\partial_{\alpha}+\frac{1}{2}\eta^{2}\partial_{\alpha\alpha}\right)h+\alpha q-\phi q^{2}\\ &\quad+\lambda^{+}\bigg{(}\max_{\delta^{+}\in\{0,1\}}\bigg{\{}\rho(\delta^{+})\mathbb{I}_{q>\underline{q}}\mathbb{E}\Big{[}\delta^{+}\frac{\Delta}{2}+h(t,\alpha+\epsilon^{+},q)\Big{]}\\ &\qquad-h(t,\alpha+\epsilon^{+},q)\Big{]}\bigg{\}}+\mathbb{E}\Big{[}h(t,\alpha+\epsilon^{+},q)-h(t,\alpha,q)\Big{]}\bigg{)}\\ &\quad+\lambda^{-}\bigg{(}\max_{\delta^{-}\in\{0,1\}}\bigg{\{}\rho(\delta^{-})\mathbb{I}_{q<\overline{q}}\mathbb{E}\Big{[}\delta^{-}\frac{\Delta}{2}+h(t,\alpha-\epsilon^{-},q)\Big{]}\\ &\qquad-h(t,\alpha-\epsilon^{-},q)\Big{]}\bigg{\}}+\mathbb{E}\Big{[}h(t,\alpha-\epsilon^{-},q)-h(t,\alpha,q)\Big{]}\bigg{)}\end{split}$ (10)
subject to the terminal condition:
@@
As previously highlighted, a major shortcoming of this model is that adverse fills can no longer be directly incorporated into the solution. This occurs because the state variable $S$ has been completely eliminated from the problem through the ansatz in Equation (9). Since assessing adverse fills requires determining whether a price change has occurred, the absence of $S$ makes this evaluation impossible. However, in Section 5, we incorporate adverse fills into our improved simulation environment, allowing us to analyze the impact of their exclusion from the model.
Lastly, by setting $\delta=0$ and assuming $\rho(\delta^{\pm})$ is either a constant or a function independent of the state variables, where for simplicity we assume that $\rho_{t}^{\delta}=\rho$ with $\rho\in(0,1)$, the optimal posting decisions can be expressed in a compact form due to the structure of the maximization terms:
$\begin{split}\delta^{+,*}(t,\alpha,q)=\mathbb{I}_{\left\{\frac{\Delta}{2}+\rho\mathbb{E}[h(t,\alpha+\epsilon^{+},q-1)-h(t,\alpha+\epsilon^{+},q)]>0\right\}\cap\{q>\underline{q}\}}\\ \delta^{-,*}(t,\alpha,q)=\mathbb{I}_{\left\{\frac{\Delta}{2}+\rho\mathbb{E}[h(t,\alpha-\epsilon^{-},q+1)-h(t,\alpha-\epsilon^{-},q)]>0\right\}\cap\{q<\overline{q}\}}.\\ \end{split}$ (12)
Here, we can see that the decision to post depends on the half-spread plus the expectation over the value functions computed at $\alpha\pm\epsilon^{\pm}$ weighted by a fill probability, where an arriving MO leads to a positive or negative jump in $\alpha$. This result is almost identical to the case in Cartea et al., (2015), except now with a non-adverse fill probability, $\rho$. This probability term can be interpreted as follows:
  • If $10<\rho<$ 1, the threshold for placing an order is more restrictive, which means that the agent is less likely to post an LO because the expected benefit could be reduced.
  • If $\rho=1$, this reduces to the original case in Cartea et al., (2015), which will be reflected in the benchmark environment. The optimal controls in Equation (12) remain the same.
@@ 5 Trading Simulation Environment
In this section, we discuss the trading simulation environment used to test the performance of the optimal MM strategy. In Cartea et al., (2015) and Jaimungal, (2019), they create a simulation environment in order to back-test this strategy, which we used as our benchmark. We will make one update to the benchmark using real LOB data, rather than simulated prices. This enabled us to back-test the strategy on real financial futures contracts data later on. We will discuss the extensions we made in our improved simulation environment, which we believe more accurately portrays how this type of MM strategy would perform in reality, albeit, still with some limitations. Specifically, we will describe how our framework tracks adverse fills, as well as an improved fill probability structure related to non-adverse fills. We will then proceed to show how this can have a major effect on the performance of the trading strategy.
@@
  • Asset price: For this purpose, we will use Futures LOB data, which was collected at the proprietary trading firm Futures First Inc. The products we will focus on include the near-term expiry contracts for ES, NQ, CL and ZN. MMs often tend to focus on the more liquid products like these, in particular for HFT, as it will enable them to trade more volume. We resampled our price data into a 1 second time interval, which was then subsequently used as the time step size in our discrete environment. The data includes up to five levels of LOB events, with each new event generating a new data point at the next timestep. When resampling the dataset into 1-second intervals, we used the standard forward fill approach which ensures continuity by propagating the last observed value into missing time slots. Since market data arrives at irregular intervals, resampling often creates gaps where no new data exists. Forward fill addresses this by carrying forward the most recent data points until an update occurs, preserving the latest market state.
@@
  • Trade order fills: The benchmark trading simulation environment assumes that all fills are non-adverse, except in the unlikely event that a fill occurs simultaneously with an adverse price movement. In our improved environment, however, we explicitly account for both adverse and non-adverse fills in the fill-tracking process. The benchmark environment also assumes that all trade orders are placed at the front of the LOB queue, meaning they are the first to be filled. In most liquid financial markets, this assumption is unrealistic. Based on this observation and empirical evidence from Tables 1 and 2, we introduced a probability function $\rho$ in Section 4 that governs the likelihood of non-adverse fills. And so, non-adverse fills can be defined as:
    $NFA_{t}$ $=\sum_{t_{i}\leq t}\delta^{+}_{t_{i}}\mathbb{I}_{\{M^{+}_{t_{i}}=1\}}\rho$ (13)
    $NFB_{t}$ $=\sum_{t_{i}\leq t}\delta^{-}_{t_{i}}\mathbb{I}_{\{M^{-}_{t_{i}}=1\}}\rho,$ (14)
    where $NFA_{t}$ and $NFB_{t}$ are counting processes for all non-adverse fills at the best ask and best bid up to and including time $t$. Here, recall $\rho$ represents the probability of an order being filled at time $t_{i}$, given that it is posted at the best bid/ask and is a non-adverse fill. Unlike the benchmark environment where $\rho=1$, in reality, $\rho$ is influenced by a number of metrics such as LOB dynamics, queue position, and MO arrival rates, where we will use a rule of thumb to account for this based on the empirical evidence shown in Section 3.2. For adverse fills, we track them as follows:
    $AFA_{t}$ $=\sum_{t_{i}\leq t}\delta_{t_{i}}^{+}\mathbb{I}_{\{AS(t_{i})<AS(t_{i+1})\}}$ (15)
    $AFB_{t}$ $=\sum_{t_{i}\leq t}\delta_{t_{i}}^{-}\mathbb{I}_{\{BS(t_{i})>BS(t_{i+1})\}}$ (16)
@@
    where $AFA_{t}$ and $AFB_{t}$ are counting processes for adverse fills at the ask and bid up to and including time $t$. Here, $BS_{t_{i}}$ and $AS_{t_{i}}$ represent the bid and ask prices. To understand this intuitively, at each point in time, the trading strategy has a LO posted on the best bid (ask) $\delta_{t_{i}}^{-}(\delta_{t_{i}}^{+})$ and the asset price at time step $t_{i+1}$ is lower (higher) than at time $t_{i}$, the traders LO was then filled at the price $BS_{t_{i}}(AS_{t_{i}})$. This is quite logical because, in reality, based on the rules of the LOB system, the asset price can not move below (above) a traders posted bid (ask) LO without filling their LO first. All other fills the trader receives will be considered non-adverse fills. To ensure consistency in tracking all fills, we combine the non-adverse and adverse fill processes as follows:
    $N_{t}^{\delta,+}$ $=\sum_{t_{i}\leq t}\left(AFA_{t_{i}}+NFA_{t_{i}}\right)$ (17)
    $N_{t}^{\delta,-}$ $=\sum_{t_{i}\leq t}\left(AFB_{t_{i}}+NFB_{t_{i}}\right)$ (18)
@@
    where $N_{t}^{\delta,+}$ and $N_{t}^{\delta,-}$ are counting processes collecting all trade order fills occurring up to and including time $t$ under this trading strategy.
@@ 6 Simulation Results
Here, we will discuss the results from testing the SOC MM strategy in our improved simulation environment and compare this with the benchmark environment. First, we would like to briefly summarize the numerical solution under our parameter set and provide a visual in Figure 3, which looks similar to the solution given in Cartea et al., (2015) and Jaimungal, (2019). We compare the solutions of the buy and sell side posts, where the top plots correspond to $\rho=1$ and the bottom plots to $\rho=0.2$, and these values were chosen as these specific solutions will be utilized in our strategy simulations in Section 5. One can see how the optimal solution varies as a function of time ($t$), short-term alpha ($\alpha$) and inventory($q$). The left (right) figure displays a visual of LOs posted on the best offer (bid), where a sell (buy) LO is posted if the MMs inventory is above (below) the surface. It is clear that a lower non-adverse fill probability, $\rho$, leads to a slightly different solution where the agent is less likely to post LOs at lower inventory levels. The top left and top right plots, as well as the bottom left and bottom right plots, appear almost identical due to the equal arrival rates of MOs on both sides of the LOB. Far from maturity, the strategy is mostly independent of time, but one can see that as the MM approaches the maturity time ($T$), the optimal strategy is essentially independent of the short-term alpha ($\alpha$), since the MM is focused solely on liquidating any remaining position.
@@
As discussed earlier, the main difference in our improved simulation environment is that adverse fills are now included in the strategy simulation process, determined according to Equation (16). Our aim in this section is to show how adverse fills and fill probabilities on non-adverse fills can significantly affect the performance of the trading strategy. We will discuss the results for CL in depth here, where the same repeated analysis for ES, NQ and ZN can be found in the appendix. Based on the empirical evidence for the distribution between adverse and non-adverse fills given in Table 2, we also set the non-adverse fill probability at $\rho=0.2$. See Table 4, below, for a list of the parameter values used. We kept most of the parameters similar to Cartea et al., (2015) and Jaimungal, (2019), however, we decreased the time step and increased the maturity time, to make the length of time more realistic to the data. Some slight adjustments to some of the other parameters were also needed to keep the numerical solution stable.
Figure 3: Optimal solution as a function of time, short-term alpha, and inventory. The top plots correspond to $\rho=1$, while the bottom plots correspond to $\rho=0.2$. Here, darker red indicates a larger long position, while darker blue indicates a larger short position.
Now, we will review the strategy simulation from our improved simulation environment alongside the performance of the benchmark environment. In Figure 4, we begin by showing a snapshot of the strategy over a random 120 second path in CL, where one can see when the MM is posted on the best bid/ask and when they receive trade order fills. The left and right figures show the strategy simulation in the benchmark and improved environments, respectively. Here, the green (blue) lines indicate when the MM is posted on the best bid (ask), the filled circles indicate when the MMs LO is filled on the best bid/ask, and the unfilled circles indicate times the MM would have been filled if they had a LO posted. One noticeable feature, in the figure on the left, is that whenever the agent is posted and the price moves through their order, they do not automatically receive a fill. This, as we mentioned earlier, is because MOs are simulated independently from the price process in the benchmark models in Cartea et al., (2015) and Jaimungal, (2019), and is contrary to what would happen in reality. And so, all the fills in this left figure would be referred to as non-adverse fills. In the right figure, one can see where all the adverse fills would have occurred (denoted by AFB and AFA). Note that this can also alter the posting strategy later on because the inventory process now evolves differently to how it would in the benchmark environment.
@@
Figure 4: A snapshot of a random strategy path in the benchmark (left) and improved (right) simulation environments. The top (black/blue), middle (black), and bottom (black/green) lines represent the best ask, midprice, and best bid, respectively. Blue (green) segments indicate the agent is posted on the best ask (bid). Closed circles show fills received, while open circles represent fills the MM would have received if posted.